OpenStudy (anonymous):

how do i start this? Find an equation for all the lines which pass through the origin and which are tangent to the parabola 푦=−푥^2+6푥−4. There may be 1, 2, or more lines.

OpenStudy (anonymous):

whoa i don't know what THAT was but it's supposed to say \[y=-x^2+6x-4\]

OpenStudy (anonymous):

start by saying that if the line is tangent to the graph then the derivative is the slope, so at that the slope between \[(0,0)\] and \[(x,-x^2+6x-4)\] must be \[-2x+6\]

OpenStudy (anonymous):

slope between \[(0,0)\ \text{ and } (x,-x^2+6x-4)\] is \[\frac{-x^2+6x-4-0}{x-0}=\frac{-x^2+6x-4}{x}\]

OpenStudy (anonymous):

set this equal to \[-2x+6\] and solve for x

OpenStudy (anonymous):

i.e solve \[\frac{-x^2+6x-4}{x}=-2x+6\] \[-x^2+6x-4=-2x^2+6x\] etc

OpenStudy (anonymous):

you have a quadratic equation, you should get two values for x

OpenStudy (amistre64):

at most 2 :)

OpenStudy (anonymous):

yes, at most!

OpenStudy (anonymous):

an in fact exactly 2, one of which is itself exactly 2 if i am not mistaken

OpenStudy (anonymous):

can you go on? i'm stuck :/ thank thus far!

OpenStudy (anonymous):

\[-x^2+6x-4=-2x^2+6x\] \[x^2-4=0\] \[(x+2)(x-2)=0\] \[x=-2, x=2\]

OpenStudy (anonymous):

thanks!